### v2

@@ Abstract
Cryptocurrencies have gained popularity across various sectors, especially in finance and investment. The popularity is partly due to their unique specifications originating from blockchain-related characteristics such as privacy, decentralisation, and untraceability. Despite their growing popularity, cryptocurrencies remain a high-risk investment due to their price volatility and uncertainty. The inherent volatility in cryptocurrency prices, coupled with internal cryptocurrency-related factors and external influential global economic factors makes predicting their prices and price movement directions challenging. Nevertheless, the knowledge obtained from predicting the direction of cryptocurrency prices can provide valuable guidance for investors in making informed investment decisions. To address this issue, this paper proposes a dynamic Bayesian network approach, which can model complex systems in multivariate settings, to predict price direction movements in the next trading day of six popular cryptocurrencies. These include Bitcoin and five altcoins (cryptocurrencies other than Bitcoin): Binance Coin, Ethereum, Litecoin, Ripple, and Tether. The efficacy of the proposed model in predicting cryptocurrency price directions is evaluated from two perspectives. Firstly, comparisons are made to five popular baseline models, namely auto-regressive integrated moving average, support vector regression, long short-term memory, random forests and support vector machines. Secondly, from a feature engineering point of view, the impact of twenty-three different features, grouped into four categories, on the DBN’s prediction performance is investigated. The experimental results demonstrate that the DBN significantly outperforms the baseline models. In addition, among the groups of features, technical indicators are found to be the most effective predictors of cryptocurrency price directions.
@@
Cryptocurrencies, Dynamic Bayesian networks, Price direction prediction, Causal feature engineering
@@ 1 Introduction
The cryptocurrency market has emerged as an important player in global financial markets despite its relatively short lifespan [1, 2]. Since the inception of cryptocurrencies in 2012, [^1] Initiated by “Bitcoin: A Peer-to-Peer Electronic Cash System” paper published by a scholar named Nakamoto Satoshi [3]. the cryptocurrency market capitalisation has soared. For instance, the total market capitalisation of cryptocurrencies witnessed an increase from 17 billion dollars at the beginning of 2017 to 1.1 trillion dollars at the beginning of 2023.222coinmarketcap.com This exponential growth has made cryptocurrencies an excellent investment opportunity and attracted a considerable number of investors to this market. Furthermore, Bitcoin is losing its dominance to altcoins, as recent studies have suggested [11, 12], highlighting the need for expanding research into altcoins along with Bitcoin. [^3] Altcoins have significantly grown in popularity; their market capitalisation surged from 9% in January 2017 to 60% in January 2023, with the number of different cryptocurrencies expanding from 50 to over 20,000, according to data from coinmarketcap.com
Despite the attractiveness of the cryptocurrency market, there are several challenges concerning the profitability of cryptocurrency trading [76]. In addition to common challenges that affect investment decisions in traditional financial assets, such as accurate price prediction and being influenced by global economic and non-economic factors [77], cryptocurrencies face distinct challenges due to their unique ecosystem. The challenges that contribute to the volatility of their prices include mining difficulty, the security of wallets and cryptocurrency exchanges, blockchain-related energy consumption, and the lack of international acceptance and legislation [4]. Moreover, the cryptocurrency market is characterised by no closed trading periods, and often impacted by public sentiment, which distinguishes it from other financial markets [5]. Cryptocurrencies exhibit higher volatility and lower liquidity than traditional financial instruments, presenting challenges for investors aiming to trade cryptocurrencies [96]. Hence, the interplay between these unique characteristics of cryptocurrencies and challenges in the financial markets adds further complexity in predicting cryptocurrency market behaviour.
Although knowledge about the future price movement directions of financial assets is useful for investors, predicting those movements is a complex task in financial time series analysis [6]. This challenge is primarily due to the inherently complex, nonlinear, and noisy attributes of financial data [7]. In addition, there are numerous factors affecting the variability in the financial data that necessitate feature engineering beyond associations derived from correlations. These factors may include market sentiment, regulatory changes, and macroeconomic indicators [91, 92, 93] Furthermore, in the context of cryptocurrency market, where multiple financial entities are represented in a multivariate time series, fluctuations in one variable can significantly impact the values of other variables [8]. Nevertheless, predicting price direction is useful for short-term investors to reduce cryptocurrency market potential risks by taking necessary steps when their investment value may increase or decrease due to market conditions [6]. Additionally, predicting the direction of prices can generate ‘buy’ and ‘sell’ signals, which can be used by algorithmic traders to develop automated trading activities [9]. Therefore, by reducing the quantitative price value prediction problem to a narrower version in terms of classifying price movement direction, it becomes possible to determine trend changes qualitatively for a given window of time [10]. Adopting this approach can potentially result in a useful technique for predicting cryptocurrency market directions and is an essential step toward designing reliable procedures for predicting the actual quantitative prices of financial assets [8].
@@
Artificial intelligence (AI), in particular, machine learning (ML) has proven to be effective in prediction across various fields, ranging from health [79, 80] to transportation [78], and finance [81]. As a result, financial technology (Fintech) [^4] The term ‘Fintech’ refers to individuals or companies that bring innovation and disruption to the financial industry by merging technological and financial capabilities [13]. companies are increasingly utilising ML techniques, such as long short-term memory [14] and artificial neural networks [15], to address the challenge of price prediction in the cryptocurrency market. However, as surveyed in [16], finding an appropriate ML technique to address the challenges of predictions with high accuracy and effective feature engineering is not straightforward. In particular, the accuracy of prediction is highly sensitive to the choice of the model, sample size and corresponding hyperparameters [17, 82, 83].
Dynamic Bayesian networks (DBNs) are versatile tools for modelling complex systems that change over time and can be used to gain knowledge about the causal relationships between variables. DBNs are a widely used technique for learning dependencies between random variables in time-series data [18]. DBNs are useful for applications such as prognosis, fault detection, reliability analysis, risk assessment, and safety evaluation [19, 94]. In particular, DBNs have been used for prediction [20, 84, 85], smoothing and filtering [21], and are frequently utilised in fields such as robotics [22], speech recognition [23], and finance [24].
@@
Our research aims at predicting the price directions of cryptocurrencies using causal feature engineering, we therefore choose DBNs for several reasons. First, DBNs efficiently describe Markov processes, requiring only the current state to determine the probability of the next state. While this construction may be challenging for a complex system such as cryptocurrency market due to the vast set of states, DBNs are incredibly useful for compactly describing combinatorial state spaces in Markov processes [95]. Moreover, DBNs establish causal relationships among all system variables, effectively addressing the challenge of understanding how fluctuations in one variable affecting cryptocurrency price movements impact other factors or the entire system [8]. [^5] We use the words ‘variables’ and ‘factors’ interchangeably throughout the paper. Compared to some ML models used for classification, which are limited to predicting only the direction of market movement without accounting for the magnitude or duration of price movement directions [5], DBNs offer the advantage of providing a probabilistic inference for both upward and downward price movements, providing an estimation of the likelihood or probability of such movements. Furthermore, with the numerous factors affecting fluctuations in financial time series data, DBNs are capable of identifying the most influential set of features within the data. Understanding the most influential factors allows for a deeper insight into the drivers affecting cryptocurrency prices, which is crucial for comprehending this unique and complex market [25]. Furthermore, the lack of explainability in some ML methods, frequently described as “black-box” models, presents challenges for their practical application. The absence of critical details, such as the importance of features, the connection between variables, and the rationale behind predictions, can hinder the adoption of ML approaches, particularly for those not familiar with ML methodologies [97]. DBNs provide a structured approach that addresses this shortcoming by offering a graphical model to represent the dynamics of systems within uncertain environments [75]. This approach is particularly beneficial for analysing the cryptocurrency market, which is characterised by high levels of volatility and complexity. Despite all these important attributes, DBNs are rarely used in the domain of cryptocurrency.
@@
In response to the challenges of price prediction for financial assets, we employ DBNs to forecast the direction of price movements in cryptocurrencies. In addition to Bitcoin, the market leader, the study investigates the effectiveness of DBNs in predicting the daily price directions of five popular altcoins, which are mainly chosen considering market capitalisation and data availability. The performance of the DBN models is compared with five frequently used baseline models, namely autoregressive integrated moving average (ARIMA), support vector regression (SVR), random forests (RF), long short-term memory (LSTM), and support vector machines (SVM) models. Moreover, the research explores whether increasing the number of features fed into the DBN leads to a monotonic improvement in precision accuracy. To test this, we investigate how different feature categories influence the performance of DBNs in predicting market directions. The contributions of our research are:
@@
  • Analysing daily price directions for six popular cryptocurrencies: Bitcoin, Binance Coin, Ethereum, Litecoin, Ripple, and Tether.
@@
  • Discovering the causal relationships among four feature groups—price information, social media activity, macro-financial data, and technical indicators—and their impact on predicting cryptocurrency prices.
  • Proposing DBNs to predict cryptocurrency prices by leveraging causal influences. Specifically, by predicting the direction of price movement, DBNs can generate probabilistic ‘buy’ and ‘sell’ signals, which can be helpful for investors in making investment decisions.
@@ 2 Related work
In this section, we review a set of recent academic publications concerning using ML models in predicting price movement directions. We first briefly consider the feature engineering side of devising ML models including incorporating technical indicators, social media data, and their impact on the accuracy of the models. We then survey the applications of BNs in predicting upward and downward trends in financial markets.
@@
Technical indicators are valuable tools for making predictions and trading decisions, offering financial market participants valuable insights into market trends. They are commonly used features in ML studies for predicting various financial markets, as seen in [27] and [28]. However, they have received less attention as input features in the cryptocurrency literature. For instance, a tree-based classification model is built by Huang et al. [29] to assess whether Bitcoin returns are predictable. They create 124 indicators based on five categories of them using Python’s TA-Lib library, which include overlap study, momentum, cycle, volatility, and pattern recognition indicators. According to their results, the model has predictive power for narrow intraday ranges of Bitcoin and outperforms the buy-and-hold strategy. The study also concludes that technical analysis is useful in the Bitcoin market, despite the fact that non-fundamental factors are the primary drivers of its value. Also, a study by Alonso-Monsalve et al. [30] predicts trends of six popular cryptocurrencies (Bitcoin, Dash, Ethereum, Litecoin, Monero, and Ripple) using eighteen technical indicators derived from one-year data. They classify one-minute trends into three categories (increase, neutral, or decrease) using four different neural network architectures, including convolutional neural networks and multilayer perceptrons. Their results suggest that all cryptocurrencies are predictable to a certain extent by using technical indicators. However, the study focuses on short-term trend prediction, which is subject to the limitations such as response times and liquidity issues. Moreover, their proposed hybrid network outperforms other models, and their performances are better at predicting Bitcoin, Ethereum, and Litecoin. In another study, Akyildirim et al. [31] test the predictability of twelve cryptocurrencies using four ML algorithms, including support vector machine (SVM) and logistic regression. For their investigations, they use past price information and eight technical indicators, including the five-day relative strength index (RSI) and simple moving average (SMA) as features for their models. The results show that all four algorithms have an average classification accuracy consistently above the 50% threshold for all cryptocurrencies.
@@
The influence of popular social platforms, such as Twitter, on cryptocurrency price movements, has received substantial attention [32, 33] in the literature. In a work by Abraham et al. [34], the authors predict changes in Bitcoin and Ethereum prices using Twitter and Google Trends data. According to their study, the volume of tweets is a predictor of price direction rather than sentiment. Moreover, they employ a sentiment analysis tool in the study and observe that the sentiment of tweets tends to remain positive, regardless of price changes in their investigation. In another study, Valencia et al. [5] use multilayer perceptron, SVM, and random forest to predict the price movement of four cryptocurrencies (Bitcoin, Ethereum, Ripple, and Litecoin) by incorporating Twitter data and raw price data as two feature groups. The Twitter data used in the study include the sentiment of daily tweets, which are grouped into four categories of positive, negative, neutral, and compound sentiments. The models’ performances are compared based on independent feature groups and their combinations. The results suggest that Twitter data alone has the potential to predict specific cryptocurrencies where the best results are obtained for Bitcoin. However, there is not a universally accepted criterion for evaluating the model performance in predicting different cryptocurrencies. For example, different models show superiority on different coins based on accuracy and precision performance criteria.
@@
Despite the considerable capacity of DBNs in modelling complex stochastic situations and detecting features with causal relationships, it has rarely been used in analysing the cryptocurrency market. However, a number of studies investigating Bayesian networks (BNs) and DBNs are presented here as examples. For instance, Wang et al. [10] apply DBNs to predict the stock market trend in both the US and China markets by incorporating nine macroeconomic factors, such as gross domestic product and monthly inflation rate, into their DBNs. The study results show that the proposed method effectively captures changes in market trends that preceded actual turning points of stocks by a lead margin of a few months. However, the models do not determine the direction of the trend changes. As another example, Jangmin et al. [35] implement a variant of DBNs to model the dynamics of trends of prices of twenty companies in the Korean stock market. The results show that the proposed model can not outperform the buy-and-hold strategy as the first baseline model, since the test period of the dataset was in its bull market, and the buy-and-hold strategy naturally wins. However, the model outperforms the triple exponential average indicator, as the second baseline model, in terms of cumulative profit. In another DBNs-related study, Wang et al. [36] investigate the potential of DBNs in predicting stock prices and generating profits in the stock market. They model the stock prices of NASDAQ and the Stock Exchange of Thailand using five years of historical intraday closing-price data. The profit generated by their model is compared with several benchmarks, including the buy-and-hold strategy, and it statistically outperforms the benchmark strategies.
Regarding studies about predicting price movement direction, to detect the upward and downward movement of stock indices, Zuo and Kita [37] utilise BNs for the task. The accuracy and total profit of the BN are compared with psychological line and trend estimation algorithms. The psychological line in the study uses a predefined formula to analyse the prices of previous days and provide insights on whether the price will increase or decrease on the next day. The proposed model demonstrates an accuracy rate of around 60%, which is approximately 10% higher than the other investment strategies investigated in the study. BNs also generate much greater profit than baseline models. In addition, Malagrino et al. [38] use BNs to investigate the impact of foreign markets on Brazil’s main stock exchange index (iBOVESPA). Two different BN topologies are designed using 24- and 48-hour timeframes, with each BN predicting the index’s next-day closing direction (up or down). The study evaluates the performance of each BN using sets of one, two, and three different indices from three continents. The results reveal that the 24-hour BNs achieve a mean accuracy rate of approximately 71%, which is better than the 68% accuracy rate achieved by its 48-hour counterpart model. Additionally, the study found that the BN models’ performance improves when fewer indices are included in the BNs.

### v3

@@ Abstract
Cryptocurrencies have gained popularity across various sectors, especially in finance and investment. Despite their growing popularity, cryptocurrencies can be a high-risk investment due to their price volatility. The inherent volatility in cryptocurrency prices, coupled with the effects of external global economic factors, makes predicting their price movements challenging. To address this challenge, we propose a dynamic Bayesian network (DBN)-based approach to uncover potential causal relationships among various features including social media data, traditional financial market factors, and technical indicators. Six popular cryptocurrencies, Bitcoin, Binance Coin, Ethereum, Litecoin, Ripple, and Tether are studied in this work. The proposed model’s performance is compared to five baseline models of auto-regressive integrated moving average, support vector regression, long short-term memory, random forests, and support vector machines. The results show that while DBN performance varies across cryptocurrencies, with some cryptocurrencies exhibiting higher predictive accuracy than others, the DBN significantly outperforms the baseline models.
@@
Cryptocurrencies , Dynamic Bayesian Networks , Price Direction Prediction , Causal Feature Engineering , Social Media , Technical Indicators
[…]
[label0]organization=School of Information Technology, Deakin University,addressline=Waurn Ponds Campus, city=Geelong, postcode=3216, state=Victoria, country=Australia
[…]
[label1]organization=Deakin Business School, Deakin University, addressline=Burwood, city=Melbourne, postcode=3125, state=Victoria, country=Australia
@@ 1 Introduction
The cryptocurrency market has emerged as an important player in global financial markets despite its relatively short lifespan [1]. Since the inception of cryptocurrencies in 2012, [^1] Initiated by the “Bitcoin: A Peer-to-Peer Electronic Cash System” paper by Nakamoto Satoshi [2]. the market capitalisation has soared, increasing from 17 billion in early 2017 to 2.6 trillion US dollars in March 2024.222coinmarketcap.com This impressive growth underscores the significance of cryptocurrencies and their impact on attracting a large number of investors.
Despite the attractiveness of the cryptocurrency market, several challenges affect the profitability of cryptocurrency trading [70]. Cryptocurrencies face common challenges encountered by traditional financial assets, such as accurate price prediction and global economic and non-economic influences [59]. These challenges are primarily due to the inherently complex, nonlinear, and noisy attributes of financial data [6] and influences of factors such as market sentiment, regulatory changes, and macroeconomic indicators [81, 82]. Additionally, cryptocurrencies encounter unique challenges due to their distinct features, which are not present in other financial markets, including mining difficulty, wallet and exchange security, blockchain energy consumption, and lack of international acceptance and regulation [3, 4, 87]. These factors contribute to higher volatility and add complexity to predicting cryptocurrency price movements [85]. Considering this complexity, providing investors with tools that can explain the relationships between these factors and predict price direction is influential. By simplifying the problem to classifying price movement direction, these tools can identify trends, improve predictive accuracy for financial assets [8, 7], and generate buy and sell signals for trading [5].
@@
Artificial intelligence (AI), particularly machine learning (ML), has proven effective in prediction across various fields [71]. Researchers and FinTechs [^3] The term “Fintech” refers to individuals or companies that bring innovation and disruption to the financial industry by merging technological and financial capabilities [11]. are increasingly utilising ML techniques, such as long short-term memory [12] and artificial neural networks [13], to address the challenge of price prediction in the cryptocurrency market. However, finding an appropriate ML technique to achieve high accuracy is complex [14]. In particular, the accuracy of prediction is sensitive to the choice of the model, sample size and corresponding hyperparameters [15, 74]. Additionally, many ML models are limited by their “black-box” nature, which lacks explainability and does not provide insights into the importance of features or the relationships between variables [86]. These limitations are particularly challenging in the context of financial markets, where understanding the underlying factors driving predictions is crucial for making informed investment decisions.
@@
To address these challenges, we use DBNs to predict the price direction of cryptocurrencies for several reasons. They offer a structured approach to modelling complex systems that change over time. They efficiently describe Markov processes, requiring only the current state to determine the probability of the next state, making them suitable for the volatile and complex cryptocurrency market [84]. Compared to some ML models used for classification, which are limited to predicting only the direction of market movement without accounting for the magnitude of price movement directions [4], DBNs offer the advantage of providing a probabilistic inference for price movement directions, estimating the magnitude and likelihood of such movements. One of the key advantages of DBNs is their explainability, which facilitates understanding the relationships among various factors. Specifically, DBNs can identify the most influential features by establishing causal relationship networks among them and offering a graphical model to represent the dynamics of systems [69]. This approach offers deeper insights into the drivers of cryptocurrency price and their impact on price fluctuations [23]. Despite all these important attributes, DBNs are rarely used in the domain of cryptocurrency.
@@
We employ DBNs to predict the direction of price movements in cryptocurrencies. In addition to Bitcoin, the market leader, the study investigates the effectiveness of DBNs in predicting the daily price directions of five popular altcoins, which are mainly chosen considering market capitalisation and data availability. The performance of the DBN models is compared with five frequently used baseline models, namely autoregressive integrated moving average (ARIMA), support vector regression (SVR), random forests (RF), long short-term memory (LSTM), and support vector machines (SVM) models. Moreover, the research explores various features to determine if increasing the number of features fed into the DBN leads to a monotonic improvement in prediction accuracy. By examining how different feature categories influence the performance of DBNs, we aim to enhance the accuracy of market direction predictions. The contributions of our research are:
  • 1.
@@
    Analysing daily price direction movements for six popular cryptocurrencies—Bitcoin, Binance Coin, Ethereum, Litecoin, Ripple, and Tether—using AI methods to evaluate predictive accuracy and provide insights into market behaviour.
  • 2.
@@
    Investigating the impact of four feature groups—price information, social media activity, macro-financial data, and technical indicators—on the accuracy of cryptocurrency price predictions, highlighting their interplay and collective influence on predictive performance.
  • 3.
    Proposing a DBN-based approach for predicting cryptocurrency prices by analysing causal relationships among market factors. This methodology generates probabilistic “buy” and “sell” signals, helping investors make informed trading decisions.
  • 4.
    Delivering explainable Bayesian networks that identify the most influential factors affecting price changes through causal inferential analysis, thereby alleviating the “black box” problem faced by many AI models.
@@ 2 Related work
In this section, we review a set of recent academic publications concerning using ML models in predicting price movement directions. We first briefly consider the feature engineering side of devising ML models including incorporating technical indicators, social media data, and their impact on the accuracy of the models. We then survey the applications of Bayesian networks (BNs) in predicting upward and downward trends in financial markets.
@@
Technical indicators are valuable tools for making predictions and trading decisions, offering financial market participants valuable insights into market trends. They are commonly used features in ML studies for predicting various financial markets, as seen in [25] and [26]. However, they have received less attention as input features in the cryptocurrency literature. For instance, Huang et al. [27] build a tree-based model to assess Bitcoin return predictability, using 124 indicators categorised into overlap study, momentum, cycle, volatility, and pattern recognition. Their model shows predictive power for Bitcoin’s narrow intraday ranges and outperforms the buy-and-hold strategy, demonstrating the utility of technical analysis in Bitcoin despite its value being driven mainly by non-fundamental factors. Another study by Alonso-Monsalve et al. [28] predicts trends for six cryptocurrencies—Bitcoin, Dash, Ethereum, Litecoin, Monero, and Ripple—using eighteen technical indicators calculated from one year of data. They classify one-minute trends into increase, neutral, or decrease categories with four neural network architectures, including convolutional neural networks and multilayer perceptrons. The study demonstrates that all cryptocurrencies can be predicted to some extent with technical indicators, with better performance for Bitcoin, Ethereum, and Litecoin. However, short-term predictions are limited by response times and liquidity issues. In another study, Akyildirim et al. [29] analysed the predictability of 12 cryptocurrencies using four ML algorithms, including support vector machine (SVM) and logistic regression. For their investigations, they use past price information and eight technical indicators, including the five-day relative strength index (RSI) and simple moving average (SMA) as features for their models. The results show that all four algorithms have an average classification accuracy consistently above the 50% threshold for all cryptocurrencies.
@@
The influence of popular social platforms, such as Twitter, [^4] The company’s name has changed to X, but we use the name that was in effect at the time of the study. on cryptocurrency price movements, has received substantial attention [30, 31] in the literature. In a work by Abraham et al. [32], the authors predict changes in Bitcoin and Ethereum prices using Twitter and Google Trends data. According to their study, the volume of tweets is a predictor of price direction rather than sentiment. Moreover, they employ a sentiment analysis tool in the study and observe that the sentiment of tweets tends to remain positive, regardless of price changes in their investigation. In another study, Valencia et al. [4] explores price movement prediction for four cryptocurrencies—Bitcoin, Ethereum, Ripple, and Litecoin—using multilayer perceptron, SVM, and random forest models. They incorporate two feature groups: raw price data and Twitter sentiment data. The Twitter data is categorised into positive, negative, neutral, and compound sentiments. The study compares the models’ performance based on independent and combined feature groups. Results indicate that Twitter sentiment data alone can predict cryptocurrency movements, with the highest accuracy achieved for Bitcoin. However, there is no universally accepted performance criterion, and different models show varying effectiveness across cryptocurrencies based on accuracy and precision.
@@
Despite the considerable capacity of DBNs in modelling complex stochastic situations and detecting features with causal relationships, it has rarely been applied in analysing the cryptocurrency market. However, studies exploring BNs and DBNs in other domains highlight their versatility and effectiveness in various applications. For instance, Wang et al. [8] use DBNs with nine macroeconomic factors to predict stock market trends in the US and China. Their method effectively detects market trend changes before actual turning points, however, it does not determine the direction of these trends, focusing instead on the timing of market shifts. Jangmin et al. [33] apply a variant of DBNs to model price trends for 20 companies in the Korean stock market. Although the model does not surpass the buy-and-hold strategy, which performs well during a bull market, it does achieve better cumulative profit compared to the triple exponential average indicator. In another study, Wang et al. [34] explore the use of DBNs for predicting stock prices and generating profits. They analyse NASDAQ and the Stock Exchange of Thailand using five years of historical intraday closing-price data. Their DBN model demonstrates significant profit generation, statistically outperforming several benchmark strategies, including the buy-and-hold approach.
To predict stock index price movements, Zuo and Kita [35] employ BNs and compare their performance against the psychological line and trend estimation algorithms. The psychological line algorithm uses historical price data to forecast future price directions. The study finds that the BN model achieves an accuracy rate of around 60%, which is approximately 10% higher than the other investment strategies considered. Additionally, BNs demonstrate significantly greater profitability than the baseline models, highlighting their effectiveness in predicting price movements. In addition, Malagrino et al. [36] use BNs to assess the influence of foreign markets on Brazil’s iBOVESPA index. They test two BN models with 24-hour and 48-hour timeframes to predict the next day’s index direction. The 24-hour BNs achieve a mean accuracy of about 71%, outperforming the 48-hour BNs, which have a 68% accuracy rate. The study also finds that including fewer indices improves BN performance.
In summary, based on the reviewed studies, while technical indicators and social media data have shown promising results, there is a notable lack of comprehensive research on their application across various cryptocurrencies. Furthermore, the potential of DBNs for cryptocurrency prediction, along with their ability to provide explainable influence networks, remains underexplored, highlighting the need for further research.